OpenStudy (anonymous):

Prove: sin θ - sin θ•cos^2 θ = sin^3 θ. You must show all work.

OpenStudy (gorv):

\[\cos^2\theta = 1-\sin^2\theta\]

OpenStudy (anonymous):

take sin theta common and then use the identity.

OpenStudy (gorv):

\[\sin \theta-\sin \theta *(1-\sin^2 \theta)\]

OpenStudy (gorv):

sin\[\sin \theta -\sin \theta +\sin^3\theta\]

OpenStudy (gorv):

=\[\sin^3 \theta\]

OpenStudy (gorv):

@?HelpMePlease?

OpenStudy (anonymous):

Thanks